OpenStudy (anonymous):

I need for someone to please explain to be the distance/time/rate formulas, and how to solve problems such as the following: Two cars leave towns 1000 km apart at the same time and travel toward each other. One car's rate is 16 km/hr more than the other's. If they meet in 5 hours, what is the rate of the faster car? I've been at it for DAYS, but I just can't get it. Please help!!!

OpenStudy (amistre64):

if we go a certain speed for a given time, we have traveled a distance. d = st

OpenStudy (amistre64):

speed of course is just a rewriting of that: speed = distance per time s = d/t

OpenStudy (amistre64):

the distance traveled by boths cars will have to add up to the total distance between them. 1000 = d1 + d2 we know that d = st, and that t=5hours soo d1 = 15 s d2 = 15 (s + 16) ... its going 16 more that the other sub them in 1000 = 15s + 15(s+16) solve for s

OpenStudy (amistre64):

pfft, 5 not 15. my fingers hate me